# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

### v1

@@ 5 Conclusions and summary
We present mid-infrared imaging and spectroscopy of the â=â7.15 Lyman-break galaxy B14-65666 taken with MIRI/JWST. The data provide the first detection of the H emission line of this source, a system of two merging galaxies at the Epoch of Reionization. The study also combines the new MIRI data with NIRCam data (Sugahara et al. 2025) and archival high angular resolution ALMA imaging (Hashimoto et al. 2018) to derive additional physical properties. The main results are summarized below:
@@
The H emitting gas shows a structure dominated by two spatially resolved sources (E and W) separated by a projected distance of 0.4 arcsec (i.e. 2.2âkpc) and relative velocity of 175âkmâs-1. The rest-frame ultraviolet light shows galaxy E as a very compact, unresolved source (upper limit for the effective radius of 63âpc) while galaxy W is more extended (348âpc) with the presence of fainter sources along an elongated structure.
@@
The SFR derived from the H luminosity is 76 and 30 Mââyr-1 (for solar metallicity as reference), for galaxies E and W respectively. Other measurements of the SFR using hybrid tracers, H and the luminosity TIR, and the UV plus TIR luminosities give values of 50 and 54 Mââyr-1, and 66 and 88 Mââyr-1 for galaxies E and W, respectively. These values are compatible within the uncertainties with the SFR derived from H. The ratio of H-to-UV based SFR (burstiness) indicates that while galaxy E is in a mild bursty phase (ratio of 1.8), galaxy W is experiencing a more continuous burst (ratio of 0.8).
@@
The equivalent width of the H line (832â±â100 and 536â±â78 , for galaxies E and W, respectively) and the ionizing photon production efficiency (â=â25.1â±â0.1 Hzâerg-1 and â=â25.5â±â0.1 Hzâerg-1) in the two galaxies are within the range of values measured in other samples of galaxies at these redshifts. Their different locations in the production efficiency âââequivalent width plane, with galaxy W above and E below the relation for high- galaxies is interpreted as due to slightly different ages in their young stellar populations.
@@
The specific star formation rate (40âââ50âGyr-1) indicates that the two galaxies involved in the B14-65666 system are experiencing an intense starburst phase, well above the typical values for galaxies of similar mass at those redshifts, and similar to less massive (â10) galaxies.
@@
The stellar mass surface density of galaxy E is extreme (â6âââMââpc-2) and close to the value measured in the nuclei of low- galaxies while galaxy W has a surface density (âMââpc-2) similar to that of high- galaxies.
@@
The H emission line profile shows a different gas kinematics, with galaxy W showing larger velocities (i.e. broader line) than galaxy E. The ionized gas kinematics traced by the H line shows differences with respect to the neutral and highly ionized gas traced by the far-infrared [CâII]â158âm and [OâIII]â88âm lines, respectively. This indicates a complex structure in the interstellar medium, likely due to the internal substructure of the young/mature stellar populations and dust distribution in these galaxies on (sub)kpc scales.
In summary, B14-65666 is a system consisting of two galaxies that are experiencing an extreme starburst phase due to the interaction/merging process. One of the galaxies (E) is already very compact and massive, with a stellar density close to that of the mature nucleus/bulge of low galaxies, while the other (W) is consistent with a more extended star formation, distributed in a few fainter clumps along a tidal tail. These results agree with recent cosmological simulations of clumpy merging systems in the early Universe (Nakazato et al. 2024). According to these simulations, this clumpy phase identified in B14-65666 is short-lived, with the bursty clumps merging in a short time scale (50 Myr) and forming a massive, single-nucleus galaxy at later times. Detailed studies of systems similar to B14-65666 will help understanding the physical conditions and mechanisms involved in the formation process of massive galaxies and their evolution in the early Universe, just a few hundred Myr after the Big Bang, and during the Epoch of Reionization.

### v2

@@ 5 Conclusions and summary
We present mid-infrared imaging and spectroscopy of the â=â7.15 Lyman-break galaxy B14-65666 taken with MIRI/JWST. The data provide the first detection of the H emission line of this source, a system of two merging galaxies at the EoR. The study also combines the new MIRI data with NIRCam data (Sugahara et al. 2025) and archival high-angular-resolution ALMA imaging (Hashimoto et al. 2018) to derive additional physical properties. The main results are summarized below.
@@
The H-emitting gas shows a structure dominated by two spatially resolved sources (E and W) separated by a projected distance of 0.4 arcsec (i.e., 2.2âkpc) and relative velocity of 175âkmâs-1. The rest-frame UV light shows galaxy E as a very compact, unresolved source (upper limit for the effective radius of 63âpc), while galaxy W is more extended (348âpc) with the presence of fainter sources along an elongated structure.
@@
The SFR derived from the H luminosity is 76 and 30 Mââyr-1 (for solar metallicity as reference), for galaxies E and W, respectively. Other measurements of the SFR using hybrid tracers; H and the luminosity TIR; and the UV-plus-TIR luminosities give values of 50 and 54 Mââyr-1 and 66 and 88 Mââyr-1 for galaxies E and W, respectively. These values are compatible within the uncertainties with the SFR derived from H. The ratio of H-to-UV-based SFR (burstiness) indicates that while galaxy E is in a mild bursty phase (ratio of 1.8), galaxy W is experiencing a more continuous burst (ratio of 0.8).
@@
The equivalent width of the H line (832â±â100 and 536â±â78 for galaxies E and W, respectively) and the ionizing photon production efficiency (â=â25.1â±â0.1 Hzâerg-1 and â=â25.5â±â0.1 Hzâerg-1) in the two galaxies are within the range of values measured in other samples of galaxies at these redshifts. Their different locations in the production efficiency equivalent width plane, with galaxy W above and galaxy E below the relation for high- galaxies is interpreted as being due to slightly different ages in their young stellar populations.
@@
The specific star formation rate (40âââ50âGyr-1) indicates that the two galaxies involved in the B14-65666 system are experiencing an intense starburst phase, well above the typical values for galaxies of similar masses at those redshifts and similar to less massive (â10) galaxies.
@@
The stellar-mass surface density of galaxy E is extreme (â6âââMââpc-2) and close to the value measured in the nuclei of low- galaxies, while galaxy W has a surface density (âMââpc-2) similar to that of a high- galaxy.
@@
The H emission-line profile shows a different gas kinematics, with galaxy W showing larger velocities (i.e., broader line) than galaxy E. The ionized gas kinematics traced by the H line shows differences with respect to the neutral and highly ionized gas traced by the far-infrared [CâII]â158âm and [OâIII]â88âm lines, respectively. This indicates a complex structure in the ISM, likely due to the internal substructure of the young and mature stellar populations and dust distribution in these galaxies on (sub)kiloparsec scales.
In summary, B14-65666 is a system consisting of two galaxies that are experiencing an extreme starburst phase due to the interaction/merging process. One of the galaxies (E) is already very compact and massive, with a stellar density close to that of the mature nucleus/bulge of low galaxies, while the other one (W) is consistent with a more extended star formation, distributed in a few fainter clumps along a tidal tail. These results agree with recent cosmological simulations of clumpy merging systems in the early Universe (Nakazato et al. 2024). According to these simulations, this clumpy phase identified in B14-65666 is short-lived, with the bursty clumps merging on a short timescale (50 Myr) and forming a massive, single-nucleus galaxy at later times. Detailed studies of systems similar to B14-65666 will help our understanding of the physical conditions and mechanisms involved in the formation process of massive galaxies and their evolution in the early Universe, just a few hundred Myr after the Big Bang and during the EoR.